OpenStudy (anonymous):

Given: f(x) = x-1/x-4 find f(1/a+1) Any Ideas guys? :?

OpenStudy (lgbasallote):

change every x with 1/a+1

OpenStudy (anonymous):

And then find the LCD which is 1/a+1 then multiply the other side of the equation to get the same denominator. After that subtract the whole number with the X?

OpenStudy (lgbasallote):

actually no...lol sorry

OpenStudy (lgbasallote):

\[\frac{x-1}{x-4} = \frac{\frac{1}{a+1} - 1}{\frac{1}{a+1} - 4}\] \[\frac{\frac{1 - (a+1)}{a+1}}{\frac{1 - 4(a+1)}{a+1}}\] got that?

OpenStudy (anonymous):

;O noope

OpenStudy (anonymous):

The LCD is a+1 right? Sorry typo

OpenStudy (lgbasallote):

okay...\[\frac{1}{a+1} - 1\] LCD is a+ 1 so we make 1 with denominator a+1 \[\frac{1}{a+1} - \frac{a+1}{a+1}\] now they are common denominatrs so just subtract the numerators copy thedenominator \[\frac{1 - (a+1)}{a+1}\] now do you get it??

OpenStudy (anonymous):

Sure did! Thanks! btw, how can I ask another question? You might now the answer.

OpenStudy (anonymous):

OR should I make another post for this question?

OpenStudy (lgbasallote):

well asking as new question would help as more people can view it

OpenStudy (anonymous):

OKay :) Thanks again!

OpenStudy (lgbasallote):

you're welcome :)